There are four newly hired salespersons at Honda Cars Taguig Branch. The

number of motorcycles

and9.sold by each newly hired salesperson last month was 4,5, 8,Complete the information needed in the table below.

a. How many samples of size 2 are possible?

b. List all possible samples of size 2.

Observation Sample


1
Expert's answer
2021-07-13T08:54:13-0400

Assuming that the samples are drawn without replacement.

a.

"N= 4 \\

n=2 \\

Number \; of \; samples = \frac{4!}{2!(4-2)!}= \frac{3 \times 4}{2}=6"

b.